OpenStudy (anonymous):

is the x intercepts of the equation f(x)=5X^2-15 (0,-15)?

OpenStudy (bahrom7893):

x intercept means y=0

OpenStudy (bahrom7893):

or f(x) = 0 0 = 5x^2-15 15=5x^2 3=x^2 x = +/-sqrt(3)

OpenStudy (bahrom7893):

so that's not an x intercept.

OpenStudy (anonymous):

ok im lost then let me show you the whole problem

OpenStudy (anonymous):

solve 5x^2=15.............thats the answer i got

OpenStudy (bahrom7893):

yes that's right.

OpenStudy (anonymous):

then find the x intercepts of f(x)=5x^2-15

OpenStudy (anonymous):

y intercept would be -15? so would the x intercept be 0?

OpenStudy (bahrom7893):

x intercept means y=0, solve for x y intercept means x=0, solve for y

OpenStudy (amistre64):

the x axis IS the y=0 axis. the x intercept IS the y=0 intecept.

OpenStudy (anonymous):

i have to type and orderd pair for the answer so i thought it would be 0,-15

OpenStudy (amistre64):

so, you have written here: when y=-15, x-0 this is NOT of the form: when y=0. x=__

OpenStudy (amistre64):

0 is not equal to -15

OpenStudy (amistre64):

f(x)= y =5x^2-15 = 0 5x^2-15 = 0 ; when x=__

OpenStudy (anonymous):

im lost lol

OpenStudy (amistre64):

5 times "something" - 15 = 0 we should know by know that: 15-15 = 0 so.. 5 times "something" has to equal 15 5 times "3" = 15 .. soo x^2 = something = 3 x^2 = 3 when x = +- sqrt(3)

OpenStudy (amistre64):

algebra-wise that is: 5x^2 - 15 = 0 +15 +15 --------------- 5x^2 = 15 /5 /5 --------------- x^2 = 3 sqrt(x^2) = +- sqrt(3) x = +- sqrt(3)

OpenStudy (amistre64):

therefore we have 2 solutions; 2, y=0 intercepts (sqrt(3),0) and (-sqrt(3),0)

OpenStudy (anonymous):

ooohh i see now lol

OpenStudy (amistre64):